OpenStudy (anonymous):

HELP PLEASE!!! Antonio has found that he can input statistical data into his computer at the rate of 2 data items faster each half hour he works. On Monday, he starts work at 9:00am, inputting at a rate of 3 data times per minute. At what rate will Antonio be inputting data into the computer by lunchtime (noon)? Show your work

OpenStudy (anonymous):

12 00 - 9 00 = 3 00 hours 3 00 hours / 0.5 hours = 6 6 * 2 = 12 12 + 3 = 15